Solve the proportion: 4/(x+1) = 3/(x+2)
Solve the proportion: x/(x+3) = (x-2)/x
Solve the proportion: 5/x = 8/15
Please i need your help ASAP! thankyou so much

Found 2 solutions by sofiyac, stanbon:
Answer by sofiyac(983)

Solve the proportion: 4/(x+1) = 3/(x+2)
4(x+2)=3(x+1)
4x+8=3x+3
x=-5
Solve the proportion: x/(x+3) = (x-2)/x
x^2=(x+3)(x-2)
x^2=x^2+3x-2x-6
x^2=x^2+x-6
0=x-6
6=x
Solve the proportion: 5/x = 8/15
5*15=8x
75=8x
9.375=x
Try to post one problem at a time next time, thanks

Answer by stanbon(75887)

Solve the proportion: 4/(x+1) = 3/(x+2)
Cross-multiply:
4(x+2) = 3(x+1)
4x+8 = 3x+3
x = -5
-------------------
Solve the proportion: x/(x+3) = (x-2)/x
Cross-multiply:
x^2 = (x+3)(x-2)
x^2 = x^2+x-6
x-6 = 0
x = 6
===================
Solve the proportion: 5/x = 8/15
x/5 = 15/8
x = 5(15/8)
x = 75/8
=============
Cheers,
Stan H.